# MemControl: Mitigating Memorization in Diffusion Models via Automated Parameter Selection

### v2

@@ Abstract
Figure 1: Conventional full fine-tuning results in the generation of nearly identical images across different seeds for the same text prompt. Memorization is evident through the replication of artifacts (red squares), which notably occurs with high precision and is consistently observed in the generated images. This replication can lead to patient information leakage and potential re-identification. Column 1 displays the original training images, while columns 2-5 show the closest generated samples across different seeds.
@@ 2.1 Memorization in Diffusion Models
Memorization in deep generative models has been explored in several contexts, including training data extraction [9, 8], content replication [51] and data copying [52]. Memorization occurs also in the medical domain and has been observed for brain MRI and chest X-ray data [1, 14], posing risks for example via patient re-identification [20]. Given the significant risks associated with memorization, several strategies have been proposed to mitigate it. Based on evidence that duplicate samples increase memorization, the most straightforward approach involves de-duplicating the training set [8, 61, 41]. However, such approach has been shown to be insufficient [52]. In text-to-image generative models, text conditioning significantly influences memorization and several mitigation strategies have been proposed based on this observation [52, 62, 44].
@@ 4.1 Experimental Setup
Experimental Settings: All experiments fine-tune a pre-trained Stable Diffusion model (v1.5) for 20,000 steps with a batch size of 128, using the AdamW optimizer [39]. The HPO search was implemented using Optuna package [2]. The FID score ($d^{fid}$) is reported on a held-out test set, and memorization, indicated by the AMD score ($d^{amd})$, is measured on the training set. The BioViL-T score is measured between the original text captions and generated images from the test set. The extraction attack is conducted between synthetic images and the training set images.
@@ 4.5 Additional Ablations
4.6 Analyzing the Best Solution
Our best solution relied on a subset of SV-DIFF learnable parameters. We analyze the mask indicating which subset of parameters should be fine-tuned. It indicates necessity of fine-tuning all of the cross-attention layers in the Stable Diffusion U-Net, which strongly aligns with the findings in [22]. Additionally, the mask suggests fine-tuning all of the self-attention layers except for those in down-blocks 2 and 3 and up-blocks 3 and 4. Adopting this strategy corresponds to fine-tuning 0.019% of the parameters in Stable Diffusion U-Net.
@@ 5 Conclusion
In this paper, we conducted the first empirical study of memorization in text-to-image diffusion models applied to medical images. We addressed the challenge of mitigating memorization from the perspective of parameter-efficient fine-tuning with capacity control. Our work demonstrated how the selection of fine-tuning parameter space influences the balance between memorization and generation quality. Through the search for optimal PEFT parameter subsets, we identified configurations that achieve a promising trade-off between generation fidelity and memorization. This introduces a novel approach to addressing the significant issue of memorization in diffusion models, distinct from the current mainstream methods based on regularization and token rewriting.

### v4

@@ Abstract
Figure 1: Conventional full fine-tuning results in the generation of nearly identical images across different seeds for the same text prompt. Memorization is evident through the replication of artefacts (red squares), which notably occurs with high precision and is consistently observed in the generated images. This replication can lead to patient information leakage and potential re-identification. Column 1 displays the original training images, while columns 2-5 show the closest generated samples across different seeds.
@@ 2.1 Memorization in Diffusion Models
Memorization in deep generative models has been explored in several contexts, including training data extraction [9, 8], content replication [51] and data copying [52]. Memorization also occurs in the medical domain and has been observed for brain MRI and chest X-ray data [1, 14], posing risks for example via patient re-identification [20]. Given the significant risks associated with memorization, several strategies have been proposed to mitigate it. Based on evidence that duplicate samples increase memorization, the most straightforward approach involves de-duplicating the training set [8, 61, 41]. However, such approach has been shown to be insufficient [52]. In text-to-image generative models, text conditioning significantly influences memorization and several mitigation strategies have been proposed based on this observation [52, 62, 44].
@@ 4.5 Additional Ablations
Analyzing the Best Solution:  Our best solution relied on a subset of SV-DIFF learnable parameters. We analyze the mask indicating which subset of parameters should be fine-tuned. It indicates necessity of fine-tuning all of the cross-attention layers in the Stable Diffusion U-Net, which strongly aligns with the findings in [22]. Additionally, the mask suggests fine-tuning all of the self-attention layers except for those in down-blocks 2 and 3 and up-blocks 3 and 4. Adopting this strategy corresponds to fine-tuning 0.019% of the parameters in Stable Diffusion U-Net.
@@ 5 Conclusion
In this paper, we conducted the first empirical study of memorization in text-to-image diffusion models applied to medical images. We addressed the challenge of mitigating memorization from the perspective of parameter-efficient fine-tuning with capacity control. Our work demonstrated how the selection of fine-tuning parameter space influences the balance between memorization and generation quality. Through the search for optimal PEFT parameter subsets, we identified configurations that achieve a promising trade-off between generation fidelity and memorization. This introduces a novel approach to addressing memorization in diffusion models, distinct from the current mainstream methods based on regularization and token rewriting.
6 Acknowledgment
Raman Dutt is supported by United Kingdom Research and Innovation (grant EP/S02431X/1), UKRI Centre for Doctoral Training in Biomedical AI at the University of Edinburgh. P. Sanchez thanks additional financial support from the School of Engineering, the University of Edinburgh. S.A. Tsaftaris acknowledges the UK’s Engineering and Physical Sciences Research Council (EPSRC) (grant EP/X017680/1) and the UKRI AI programme and EPSRC, for CHAI - EPSRC AI Hub for Causality in Healthcare AI with Real Data (grant EP/Y028856/1).